# Uncovering flow and deformation regimes in the coupled fluid-solid vestibular system

### v1

@@ 3.2.1 Sample head rotation
To visualize the different regimes we have identified above, we solve the equation for the pressure jump (24a) with a specific choice for the forcing $f(t)$. While many forms of $f(t)$ could be considered, we are motivated by a clinical head manoeuvre that may be thought of as modelling a slow rotation of the head from right to left. Although other clinical models are described by high-order polynomials (Boselli et al., 2009, 2013), we choose a particular form that facilitates an analytical solution of the equations, namely:
$f(t)=\frac{\mathrm{d}F}{\mathrm{d}t}=\begin{cases}\sin 2\pi t&t\in(0,1)\\ 0&t>1.\end{cases}$ (34)
@@
$-\Delta p=\beta\int_{0}^{t}f(\tau)e^{-\gamma\kappa(t-\tau)}\mathrm{d}\tau=\begin{cases}\beta\frac{2\pi e^{-\gamma\kappa t}+\gamma\kappa\sin 2\pi t-2\pi\cos 2\pi t}{\gamma^{2}\kappa^{2}+4\pi^{2}}&t<1,\\ \frac{2\pi\beta}{4\pi^{2}+\gamma^{2}\kappa^{2}}e^{-\gamma\kappa t}(1-e^{-\gamma\kappa})&t>1.\end{cases}$ (35)
Here, $\beta=\pi/[8I_{4}\alpha(b)]$ and $\gamma=\kappa/[16\alpha(b)I_{4}]$. The cupular deformation, which in this regime is proportional to the pressure jump, is plotted in the left panel of Figure 5 for different values of $\kappa$, from which we can clearly see the transition from $\Delta p$ tracking the angular velocity $F(t)$ for small $\kappa$ to tracking the angular acceleration $f(t)$ for large $\kappa$, as expected from the preceding analysis. There is an interesting transition region for $\kappa\sim 1$, where we can see an “overshoot” region at the end of the manoeuvre that has not decayed. This is not the case for either of the limiting regions, where the pressure jump (and cupular displacement) is identically zero after the completion of the head turn.
@@ 4 Numerical simulations
To test the validity of our asymptotic approach, we return to the numerical simulations in COMSOL as presented in §2.2 but now imposing within the numerical scheme the forcing given by (34), and varying the Young’s modulus of the solid material to change $\kappa$. We perform two direct comparisons, appearing in Figures 6 and 7. Figure 6 plots the cupular pressure jump $\Delta p$ as a function of time, while Figure 7 plots the axial velocity profiles across the cross section, sampled at several time points for different values of $\kappa$ in a region of the canal far from the cupula. In both figures, we observe excellent agreement between theory and numerics. In particular, the breaking of symmetry in the flow profile for large $\kappa$ is easily observed in Figure 7, and the theoretical profile captures the trend and profile shape very well.
@@
We remark that although the symmetry breaking might suggest a breakdown of the asymptotic order, with the second term dominating the first in the series, subsequent terms are well behaved and the series is not divergent. This may be inferred from the agreement between the model solution and the numerical solution to the full nonlinear problem, even for large $\kappa$ values where the symmetry is broken. Moreover, as the symmetry breaking occurs because the leading order term shrinks and the correction retains its size (rather than growing), we expect higher order terms to retain their sizes too, preserving the asymptotic order of the solution.
We have thus far considered slow movements, so that $\mathcal{T}>1\mathrm{~{}s}$ and $\mathrm{St}\ll 1$. However, for faster movements, typically when $\mathcal{T}<1\mathrm{~{}s}$, the Stokes number is no longer negligible and inertial terms must be retained in the analysis (see §2.4). However, the Stokes number may also become non-negligible for other reasons — for example, some authors report slightly thicker semicircular canals ($a$ slightly larger), so that the Stokes number is considerably larger than expected due to its quadratic dependence on velocity. To this end, in the next section we consider flows with a finite Stokes number.
@@
Figure 7: The numerically obtained velocity profiles (blue markers) and theoretical predictions (black solid curves) from (20a) and (20b). As $\kappa$ increases, the velocity profile ceases to be symmetric around $\kappa\approx 10^{3}$. The imposed rotation is given by (34), sampled at 7 different times. Parameter values are the same as in fig. 6.
@@ 5 Effect of fluid inertia
While considering the inertialess limit of $\mathrm{St}\ll 1$ facilitated analytical progress, there are several circumstances in which inertia may become important, e.g. faster movements or larger canals. Therefore, we consider the effect of fluid inertia by retaining the $\mathcal{O}(\mathrm{St})$ terms in the governing equations (17a). Proceeding as in the previous section and seeking a solution in powers of the small aspect ratio $\epsilon$, we seek a solution for the leading order axial velocity of the form
@@
Here, $\phi_{n}$ are the eigenfunctions for the Laplacian in a cylinder of local radius $a(s)$, subject to Dirichlet boundary conditions (Batchelor, 1973); the $\lambda_{n}$ are the zeros of the Bessel function of the first kind $J_{0}(z)$, thereby ensuring that the axial velocity satisfies the no-slip condition at $r=a(s)$ in the rotating frame. Thus, substituting (37) into the momentum equation, and using the orthogonality properties of Bessel functions we find
$\operatorname{St}\frac{\partial c_{n}}{\partial t}=-\frac{2}{\lambda_{n}J_{1}(\lambda_{n})}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)-\frac{\lambda_{n}^{2}}{a(s)^{2}}c_{n},$ (38a)
$c_{n}(t,s)=-\frac{2}{\lambda_{n}J_{1}(\lambda_{n})}\int_{0}^{t}\left(f(\tau)+\kappa\frac{\partial p_{0}}{\partial s}\right)\mathcal{K}_{n}(t-\tau,s;\mathrm{St})~{}\mathrm{d}\tau,$ (38b)
@@
$\begin{split}Q_{0}&=2\pi\int_{0}^{a(s)}rw_{0}(r,s,t)~{}\mathrm{d}r=2\pi a(s)^{2}\int_{0}^{1}\rho w_{0}(a(s)\rho,s,t)~{}\mathrm{d}\rho\\ &=-4\pi a(s)^{2}\sum_{n=1}^{\infty}\left[\lambda_{n}^{-2}\int_{0}^{t}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)\mathcal{K}_{n}(t-\tau,s;\mathrm{St})~{}\mathrm{d}\tau\right].\end{split}$ (39)
@@
$\int_{0}^{a(0)}r\frac{\partial\eta_{0}}{\partial t}\mathrm{d}r=-2a^{2}\frac{1}{\operatorname{St}}\sum_{n=1}^{\infty}\left[\frac{e^{-\lambda_{n}^{2}t/(a^{2}\operatorname{St})}}{\lambda_{n}^{2}}\int_{0}^{t}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)e^{\lambda_{n}^{2}\tau/(a^{2}\operatorname{St})}~{}\mathrm{d}\tau\right],$ (40a)
$\nabla^{4}\eta_{0}-\nabla^{2}\eta_{0}+\varrho\left(\mathrm{St}\frac{\partial^{2}\eta_{0}}{\partial t^{2}}+f(t)\right)=\Delta p_{0}=\int_{0}^{2\pi}\frac{\partial p_{0}}{\partial s}~{}\mathrm{d}s.$ (40b)
For an arbitrary inner radius $a(s)$, (40) can be solved using the Laplace transform, as shown in Appendix C. Before tackling the general case, we focus on the simple case where the tube radius is uniform, $a(s)\equiv 1$.
@@ 5.1 A simple example
$\int_{0}^{1}\frac{\partial\eta_{0}}{\partial t}r\mathrm{d}r=-2\sum_{n=1}^{\infty}\left[\lambda_{n}^{-2}\int_{0}^{t}\left(f(\tau)+\kappa\frac{\Delta p_{0}}{2\pi}\right)\mathcal{K}_{n}(t-\tau;\mathrm{St})~{}\mathrm{d}\tau\right].$ (41)
To transform this equation into a more manageable form, we define a complete kernel $\mathcal{K}(x;\mathrm{St})=\mathrm{St}^{-1}\sum_{n=1}^{\infty}\lambda_{n}^{-2}e^{-\lambda_{n}^{2}x/\mathrm{St}}$. Finally, we may directly substitute $\Delta p_{0}$ from (40b) into (39) to obtain a single equation for the cupular deflection $\eta_{0}(r,t)$:
$\int_{0}^{1}\frac{\partial\eta_{0}}{\partial t}r~{}\mathrm{d}r=-2\bar{f}(t)-\frac{\kappa}{\pi}\int_{0}^{t}\left[\nabla^{4}\eta_{0}-b^{-1}\nabla^{2}\eta_{0}+\varrho\left(\mathrm{St}\frac{\partial^{2}\eta_{0}}{\partial t^{2}}+f(\tau)\right)\right]\mathcal{K}(t-\tau;\mathrm{St})~{}\mathrm{d}\tau,$ (42)
where we have introduced $\bar{f}(t)=\int_{0}^{t}f(\tau)\mathcal{K}(t-\tau;\mathrm{St})~{}\mathrm{d}\tau$.
Equation (42) is an integro-PDE for $\eta$ with an exponential-like kernel. Seeking a separable solution $\eta_{0}=\bar{\eta}(r)\Delta p_{0}$ when $\varrho\ll 1$ will lead to an integro-differential equation for $\Delta p_{0}(t)$
$\alpha(b)\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-2\bar{f}-\frac{\kappa}{\pi}\int_{0}^{t}\Delta p_{0}(\tau)\mathcal{K}(t-\tau;\mathrm{St})\mathrm{d}\tau.$ (43)
Here, $\bar{\eta}$ and $\alpha(b)$ are the same as in the $\mathrm{St}\ll 1$ limit, given in (25). Equation (43) may be efficiently solved numerically by truncating the infinite series in the kernel and transforming the integral equation into a system of ODEs. This is a standard calculation, with details given in Appendix B.
@@ 5.2 Fluid inertia can make the cupula underdamped
To understand the effect of inertia in the cupular response, we first consider the case of a cupula that is initially stretched by a pressure jump $\Delta p_{0}(t=0)=1$ in a frame rotating at constant speed, so that $f(t)=0$.
@@
Figure 8: Left panel: Solution to (43), when $f(t)=0$ and $\Delta p_{0}(t=0)=1$ for different values of the Stokes number, showing underdamped dynamics for large enough $\mathrm{St}$. Central panel: Bifurcation diagram, showing the evolution of $\mathrm{Re}(\bar{\omega})$ (blue) and $\mathrm{Im}(\bar{\omega})$ (red). Markers represent the numerically obtained solution from (44) and dashed lines the analytical approximation (45). Right panel: bifurcation diagram for $f(t)\sim e^{2\pi it}$, showing how the transition between the two regimes $\Delta p_{0}\sim F(t)$ and $\Delta p_{0}\sim f(t)$ depends on both $\kappa$ and $\mathrm{St}$. Colour represents the complex angle of $\chi$, with blue $\arg(\chi)=\pi/2$ and red representing $\arg(\chi)=0$.
To understand this change in behaviour as $\mathrm{St}$ is increased, we seek an exponential ansatz to solve (43), with $\Delta p_{0}\sim e^{-\omega t}$ in the simple case when $f(t)=0$. Direct substitution yields the following condition:
$\frac{\alpha(b)\pi}{\kappa\mathrm{St}}\bar{\omega}=\sum_{n=0}^{\infty}\frac{1}{\lambda_{n}^{2}}\frac{1}{\lambda_{n}^{2}-\bar{\omega}},$ (44)
where the rescaled growth rate is $\bar{\omega}=\mathrm{St}\,\omega$. Equation (44) provides an equation for $\bar{\omega}$ depending on the parameter $q=\alpha\pi/(\kappa\mathrm{St})$. Analytical progress can be made by truncating the sum at $n=0$, i.e. considering only the first term. This leads to
$\bar{\omega}=\frac{\lambda_{0}^{2}}{2}\left[1\pm\left(1-\frac{4}{q\lambda_{0}^{6}}\right)^{1/2}\right].$ (45)
From this, we can identify the critical value for the parameter $q$ at which the transition to underdamped dynamics occurs: $q_{c}=4/\lambda_{0}^{6}$.
A natural question to ask now is if fluid inertia can alter the development of the two flow regimes outlined previously in Section 3.2? Proceeding as before, we assume a forcing $f(t)=Be^{2\pi it}$, and try an ansatz $\Delta p_{0}=Ae^{2\pi it}$. Substitution into the integral equation (43) and neglecting contributions from the initial conditions yields
$2\pi i\alpha A=\sum_{n=0}^{\infty}\frac{2B+A\pi\kappa}{\lambda_{n}^{2}(2\pi i\mathrm{St}+\lambda_{n}^{2})}$ (46)
It is convenient to define the function $\mathcal{G}:\mathbb{R}\rightarrow\mathbb{C}$, given by $\mathcal{G}(\mathrm{St})=\sum_{n=0}^{\infty}\lambda_{n}^{-2}/(2\pi i\mathrm{St}+\lambda_{n}^{2})$. We find that the response (characterized by $A$) is related to the forcing (characterized by $B$) through
$A=\frac{-2B\mathcal{G}(\mathrm{St})}{2\pi i\alpha+\pi\kappa\mathcal{G}(\mathrm{St})}.$ (47)
@@
$\chi=\frac{\mathcal{G}(\mathrm{St})}{2\pi i\alpha+\pi\kappa\mathcal{G}(\mathrm{St})}$ (48)
will determine if the deformation follows the angular velocity (if the angle is close to $\pi/2$) or the angular acceleration (when the angle is close to 00 or multiples of $\pi$). To achieve analytical progress we truncate the sum in $\mathcal{G}$, keeping only the first term, and we find
$\mathrm{Re}(\chi)=\pi\kappa-4\pi^{2}\mathrm{St}\alpha\lambda_{0}^{2},\quad\mathrm{Im}(\chi)=2\pi\alpha\lambda_{0}^{4}.$ (49)
@@
$\left|\frac{\kappa}{2\lambda_{0}^{4}\alpha}-\frac{2\pi\mathrm{St}}{\lambda_{0}^{2}}\right|=1.$ (50)
In Figure 8(c), we plot the argument of $\chi$ as a function of $\kappa$ and $\mathrm{St}$, indicating as well the approximate bifurcation curves given by (50). This diagram indicates where in the $\kappa$-$\mathrm{St}$ phase space the cupula deflection follows the angular velocity versus the angular acceleration. For small values of the Stokes number $\mathrm{St}$ we recover the previous picture, where $\kappa\ll 1$ indicates the deformation follows the angular velocity and $\kappa\gg 1$ indicates the response is guided by the angular acceleration. But we also find a nonlinear dependence on $\mathrm{St}$ for non-small Stokes numbers. For given $\kappa$ less than about 100, a transition to angular acceleration tracking occurs for $\mathrm{St}$ greater than about 1, meaning that the cupula system may be tuned to follow angular acceleration even for small cupula stiffness if the Stokes number is high enough. This transition point increases for larger $\kappa$, as indicated by the blue wedge region in Figure 8(c), meaning that an orders of magnitude higher $\kappa$ is possible with cupular deflection still tracking angular velocity, if the Stokes number is accordingly increased in a very particular way.
We may interpret the effect of high fluid inertia by considering the response of the cupula as the forcing frequency is increased. For small forcing frequencies the deformation will follow the angular acceleration, and as the forcing frequency is increased the cupula will start deforming in phase with the angular velocity, as expected from § 3.2; this is well known in the vestibular literature (Benson, 1990). However, our results from this section suggest that when the forcing frequency is further increased (in humans, to about 100 Hz), the cupular deformation will again be in phase with the angular acceleration.

### v2

@@ 3.2.1 Sample head rotation
To visualize the different regimes we have identified above, we solve the equation for the pressure jump (31a) with a specific choice for the forcing $\dot{\Omega}(t)$. While many forms of $\dot{\Omega}(t)$ could be considered, we are motivated by a clinical head manoeuvre that may be thought of as modelling a slow rotation of the head from right to left. Although other clinical models are described by high-order polynomials (Boselli et al., 2009, 2013), we choose a particular form that facilitates an analytical solution of the equations, namely:
$\dot{\Omega}(t)=\begin{cases}\sin 2\pi t&t\in(0,1)\\ 0&t>1.\end{cases}$ (39)
@@
$\Delta p_{0}=2\pi\gamma\kappa\int_{0}^{t}\Omega(\tau)e^{-\gamma\kappa(t-\tau)}\mathrm{d}\tau=\begin{cases}2\pi\gamma\kappa\frac{2\pi e^{-\gamma\kappa t}+\gamma\kappa\sin 2\pi t-2\pi\cos 2\pi t}{\gamma^{2}\kappa^{2}+4\pi^{2}}&t<1,\\ \frac{4\pi^{2}\gamma\kappa}{4\pi^{2}+\gamma^{2}\kappa^{2}}e^{-\gamma\kappa t}(1-e^{-\gamma\kappa})&t>1.\end{cases}$ (40)
Here, $\gamma=1/[16\alpha_{0}(\beta)I_{4}]$ accounts for domain irregularities and cupular thickness. The cupular deformation, which in this regime is proportional to the pressure jump, is plotted in the left panel of Figure 5 for different values of $\kappa$, from which we can clearly see the transition from $\Delta p$ tracking the angular velocity $\Omega(t)$ for small $\kappa$ to tracking the angular acceleration $\dot{\Omega}(t)$ for large $\kappa$, as expected from the preceding analysis. There is an interesting transition region for $\kappa\sim 1$, where we can see an “overshoot” region at the end of the manoeuvre that has not decayed. This is not the case for either of the limiting regions, where the pressure jump (and cupular displacement) is identically zero after the completion of the head turn.
@@ 4 Numerical simulations
To test the validity of our asymptotic approach, we return to the numerical simulations in COMSOL as presented in §2.2 but now imposing within the numerical scheme the forcing given by (39), and varying the Young’s modulus of the solid material to change $\kappa$. We perform two direct comparisons, appearing in Figures 6 and 7. Figure 6 plots the cupular pressure jump $\Delta p$ as a function of time, while Figure 7 plots the axial velocity profiles across the cross section, sampled at several time points for different values of $\kappa$ in a region of the canal far from the cupula ($s=\pi$). In both figures, we observe excellent agreement between theory and numerics. In particular, the breaking of symmetry in the flow profile for large $\kappa$ is easily observed in Figure 7, and the theoretical profile captures the trend and profile shape very well.
@@
We remark that although the symmetry breaking might suggest a breakdown of the asymptotic order, with the second term dominating the first in the series, subsequent terms are well behaved and the series is not divergent. It can be seen from the equations that the $\mathcal{O}(\epsilon^{n})$ equations will be of the same form as (16a) but with forcing terms depending on the $\mathcal{O}(\epsilon^{m})$ solutions, with $m<n$ and no division by quantities shrinking to zero as $\epsilon\rightarrow 0$. This may be inferred from the agreement between the model solution and the numerical solution to the full nonlinear problem, even for large $\kappa$ values where the symmetry is broken. Moreover, as the symmetry breaking occurs because the leading order term shrinks and the correction retains its size (rather than growing), we expect higher order terms to retain their sizes too, preserving the asymptotic order of the solution.
We have thus far considered slow movements, so that $\mathcal{T}>1\mathrm{~s}$ and $\mathrm{St}\ll 1$. However, for faster movements, typically when $\mathcal{T}<1\mathrm{~s}$, the Stokes number is no longer negligible and inertial terms must be retained in the analysis (see §2.4). However, the Stokes number may also become non-negligible for other reasons — for example, some authors report slightly thicker semicircular canals ($a$ slightly larger), so that the Stokes number is considerably larger than expected due to its quadratic dependence on radius. To this end, in the next section we consider flows with a finite Stokes number.
@@
Figure 7: The numerically obtained velocity profiles (blue markers) and theoretical predictions (black solid curves) from (23a) and (23b), sampled at $s=\pi$ (the furthest location from the cupula). As $\kappa$ increases, the velocity profile ceases to be symmetric around $\kappa\approx 10^{3}$. The imposed rotation is given by (39), with the simulation output sampled at 7 different times. Parameter values are the same as in figure 6.
@@ 5 Effect of fluid inertia
While considering the inertialess limit of $\mathrm{St}\ll 1$ facilitated analytical progress, there are several circumstances in which inertia may become important, e.g. faster movements or larger canals. Therefore, we consider the effect of fluid inertia by retaining the $\mathcal{O}(\mathrm{St})$ terms in the governing equations (16a). Proceeding as in the previous section and seeking a Fourier-Bessel series solution for the leading order axial velocity of the form
@@
Here, $\phi_{n}$ are the eigenfunctions for the Laplacian in a cylinder of local radius $a(s)$, subject to Dirichlet boundary conditions (Batchelor, 1973); the $\lambda_{n}$ are the zeros of the Bessel function of the first kind $J_{0}(z)$, thereby ensuring that the axial velocity satisfies the no-slip condition at $r=a(s)$ in the rotating frame. Thus, substituting (42) into the momentum equation (10c), keeping only the leading order terms, and using the orthogonality properties of Bessel functions we find
$\frac{1}{r}\frac{\partial}{\partial r}$ $\left(r\frac{\partial\phi_{n}}{\partial r}\right)=-\frac{\lambda_{n}^{2}}{a(s)^{2}}\phi_{n}$ (43a)
$\int_{0}^{a(s)}r\phi_{m}(r,s)\phi_{n}(r,s)\mathrm{d}r$ $=\frac{a(s)^{2}}{2}\delta_{mn}J_{1}(\lambda_{n})^{2},\quad\int_{0}^{a(s)}r\phi_{n}(r,s)\mathrm{d}r=\frac{a(s)^{2}J_{1}(\lambda_{n})}{\lambda_{n}},$ (43b)
$\mathrm{St}\frac{\partial c_{n}}{\partial t}$ $=-\frac{2}{\lambda_{n}J_{1}(\lambda_{n})}\left(\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}\right)-\frac{\lambda_{n}^{2}}{a(s)^{2}}c_{n},$ (43c)
$c_{n}(t,s)$ $=-\frac{2}{\lambda_{n}J_{1}(\lambda_{n})}\int_{0}^{t}\left(\dot{\Omega}(\tau)+\frac{\partial p_{0}}{\partial s}\right)\mathcal{K}_{n}(t-\tau,s;\mathrm{St})~\mathrm{d}\tau,$ (43d)
@@
$\begin{split}Q_{0}&=2\pi\int_{0}^{a(s)}rw_{0}(r,s,t)~\mathrm{d}r=2\pi a(s)^{2}\int_{0}^{1}\rho w_{0}(a(s)\rho,s,t)~\mathrm{d}\rho\\ &=-4\pi a(s)^{2}\sum_{n=1}^{\infty}\left[\lambda_{n}^{-2}\int_{0}^{t}\left(\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}\right)\mathcal{K}_{n}(t-\tau,s;\mathrm{St})~\mathrm{d}\tau\right].\end{split}$ (44)
@@
$\int_{0}^{a(0)}r\frac{\partial\eta_{0}}{\partial t}\mathrm{d}r=-2a^{2}\frac{1}{\operatorname{St}}\sum_{n=1}^{\infty}\left[\frac{1}{\lambda_{n}^{2}}\int_{0}^{t}\left(\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}\right)e^{-\lambda_{n}^{2}(\tau-t)/(a^{2}\operatorname{St})}~\mathrm{d}\tau\right],$ (45a)
$\eta_{0}(r,t)=\Delta p_{0}\frac{\bar{\eta}_{0}(r)}{\kappa},\quad\Delta p_{0}=\int_{0}^{2\pi}\frac{\partial p_{0}}{\partial s}~\mathrm{d}s.$ (45b)
Therefore, we may write this as a single equation for the pressure gradient
$\frac{\alpha_{0}(\beta)}{\kappa}\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-2a^{2}\frac{1}{\operatorname{St}}\sum_{n=1}^{\infty}\left[\frac{1}{\lambda_{n}^{2}}\int_{0}^{t}\left(\dot{\Omega}(t)+\frac{\partial p_{0}}{\partial s}\right)e^{-\lambda_{n}^{2}(\tau-t)/(a^{2}\operatorname{St})}~\mathrm{d}\tau\right],$ (46)
where $\alpha_{0}(\beta)$ is given in (30). For an arbitrary inner radius $a(s)$, (46) can be solved using the Laplace transform, as shown in Appendix D. Before tackling the general case, we focus on the simple case where the tube radius is uniform, $a(s)\equiv 1$.
@@ 5.1 A simple example
$\frac{\alpha_{0}(\beta)}{\kappa}\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-2\sum_{n=1}^{\infty}\left[\lambda_{n}^{-2}\int_{0}^{t}\left(\dot{\Omega}(\tau)+\frac{\Delta p_{0}}{2\pi}\right)\mathcal{K}_{n}(t-\tau;\mathrm{St})~\mathrm{d}\tau\right].$ (47)
To transform this equation into a more manageable form, we define a complete kernel $\mathcal{K}(x;\mathrm{St})=\mathrm{St}^{-1}\sum_{n=1}^{\infty}\lambda_{n}^{-2}e^{-\lambda_{n}^{2}x/\mathrm{St}}$.
$\frac{\alpha_{0}(\beta)}{\kappa}\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-2\bar{\dot{\Omega}}(t)-\frac{1}{\pi}\int_{0}^{t}\Delta p_{0}(\tau)\mathcal{K}(t-\tau;\mathrm{St})\mathrm{d}\tau.$ (48)
where we have introduced $\bar{\dot{\Omega}}(t)=\int_{0}^{t}\dot{\Omega}(\tau)\mathcal{K}(t-\tau;\mathrm{St})~\mathrm{d}\tau$. Equation (48) may be efficiently solved numerically by truncating the infinite series in the kernel and transforming the integral equation into a system of ODEs. This is a standard calculation, with details given in Appendix C.
@@ 5.2 Fluid inertia can make the cupula underdamped
To understand the effect of inertia in the cupular response, we first consider the case of a cupula that is initially stretched by a pressure jump $\Delta p_{0}(t=0)=1$ in a frame rotating at constant speed, so that $\Omega(t)=0$.
@@
[…]
Figure 8: Left panel: Solution to (48), when $\dot{\Omega}(t)=0$ and $\Delta p_{0}(t=0)=1$ for different values of the Stokes number, showing underdamped dynamics for large enough $\mathrm{St}$. Central panel: Bifurcation diagram, showing the evolution of $\mathrm{Re}(\bar{\omega})$ (blue) and $\mathrm{Im}(\bar{\omega})$ (red). Markers represent the numerically obtained solution from (49) and dashed lines the analytical approximation (50). Right panel: bifurcation diagram for $\dot{\Omega}(t)\sim e^{it}$, showing how the transition between the three regimes depends on both $\kappa$ and $\mathrm{St}$. Colour represents the complex angle of $\chi$, with purple representing $\arg(\chi)=0$, green is $\arg(\chi)=\pi/2$ and yellow is $\arg(\chi)=\pi$.
To understand this change in behaviour as $\mathrm{St}$ is increased, we seek an exponential ansatz to solve (48), with $\Delta p_{0}\sim e^{-\omega t}$ in the simple case when $\dot{\Omega}(t)=0$. Direct substitution yields the following condition:
$\frac{\alpha_{0}(\beta)\pi}{\kappa\mathrm{St}}\bar{\omega}=\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}^{2}}\frac{1}{\lambda_{n}^{2}-\bar{\omega}},$ (49)
where the rescaled growth rate is $\bar{\omega}=\mathrm{St}\,\omega$. Equation (49) provides an equation for $\bar{\omega}$ depending on the parameter $q=\alpha_{0}(\beta)\pi/(\kappa\mathrm{St})=\alpha_{0}(\beta)\pi\rho\nu^{2}R/(Ea^{3})$, which is independent of the forcing timescale $\mathcal{T}$. Analytical progress can be made by truncating the sum at $n=1$, i.e. considering only the first term. This leads to
$\bar{\omega}=\frac{\lambda_{1}^{2}}{2}\left[1\pm\left(1-\frac{4}{q\lambda_{1}^{6}}\right)^{1/2}\right].$ (50)
From this, we can identify the critical value for the parameter $q$ at which the transition to underdamped dynamics occurs: $q_{c}=4/(\lambda_{1})^{6}\approx 0.0207$.
A natural question to ask now is if fluid inertia can alter the development of the two flow regimes outlined previously in Section 3.2? Proceeding as before, we assume a forcing $\dot{\Omega}(t)=Be^{it}$, and try an ansatz $\Delta p_{0}=Ae^{it}$. Substitution into the integral equation (48) and neglecting contributions from the initial conditions yields
$\frac{i\alpha_{0}(\beta)A}{\kappa}=-\sum_{n=1}^{\infty}\frac{2B+A/\pi}{\lambda_{n}^{2}(i\mathrm{St}+\lambda_{n}^{2})}$ (51)
It is convenient to define the function $\mathcal{G}:\mathbb{R}\rightarrow\mathbb{C}$, given by $\mathcal{G}(\mathrm{St})=\sum_{n=1}^{\infty}\lambda_{n}^{-2}/(i\mathrm{St}+\lambda_{n}^{2})$. We find that the response (characterized by $A$) is related to the forcing (characterized by $B$) through
$A=\frac{-2B\mathcal{G}(\mathrm{St})}{i\alpha_{0}(\beta)/\kappa+\mathcal{G}(\mathrm{St})/\pi}.$ (52)
@@
$\chi=\frac{-\mathcal{G}(\mathrm{St})}{\pi i\alpha_{0}(\beta)+\kappa\mathcal{G}(\mathrm{St})}=\frac{-1}{\kappa+\pi i\alpha_{0}(\beta)/\mathcal{G}(\mathrm{St})}$ (53)
will determine if the deformation follows the angular velocity (if the angle is close to $\pi/2$) or the angular acceleration (when the angle is close to 0 or multiples of $\pi$). To achieve analytical progress we truncate the sum in $\mathcal{G}$, keeping only the first term, and we find
$\frac{\mathrm{Re}(\chi)}{\mathrm{Im}(\chi)}=-\frac{\kappa-\pi\alpha_{0}(\beta)\lambda_{1}^{2}\mathrm{St}}{\pi\lambda_{1}^{4}\alpha_{0}(\beta)}.$ (54)
@@
$\left|\frac{\kappa}{\pi\lambda_{1}^{4}\alpha_{0}(\beta)}-\frac{\mathrm{St}}{\lambda_{1}^{2}}\right|=1.$ (55)
In figure 8(c), we plot the argument of $\chi$ as a function of $\kappa$ and $\mathrm{St}$, indicating as well the approximate bifurcation curves given by (55). This diagram indicates where in the $\kappa$-$\mathrm{St}$ phase space the cupula deflection follows the angular acceleration, angular velocity or the angular displacement. For small values of the Stokes number $\mathrm{St}$ we recover the previous picture, where $\kappa\ll 1$ indicates the deformation follows the angular velocity ($\mathrm{arg}(\chi)=\pi/2$) and $\kappa\gg 1$ indicates the response is guided by the angular acceleration ($\mathrm{arg}(\chi)=\pi$). But we also find a dependence on $\mathrm{St}$ for non-small Stokes numbers. For given $\kappa$ less than about 100, a transition to angular displacement tracking occurs for $\mathrm{St}$ greater than about 1, meaning that the cupula system may be tuned to follow angular displacement even for small cupula stiffness if the Stokes number is high enough (characterized by $\mathrm{arg}(\chi)=0$ in figure 8(c)). This transition point increases for larger $\kappa$, as indicated by the blue wedge region in figure 8(c), meaning that an orders of magnitude higher $\kappa$ is possible with cupular deflection still tracking angular velocity, if the Stokes number is accordingly increased in a very particular way. In the new regime for $\mathrm{St}\gg 1$ (and moderate $\kappa$) we find the pressure jump is approximately:
$\Delta p_{0}(t)=-\frac{\kappa}{2\alpha_{0}(\beta)\mathrm{St}}\int_{0}^{t}\Omega(\tau)\mathrm{d}\tau,$ (56)
and indeed proportional to the angular displacement.
We may interpret the effect of high fluid inertia by considering the response of the cupula as the forcing frequency is increased. For small forcing frequencies the deformation will follow the angular acceleration, and as the forcing frequency is increased the cupula will start deforming in phase with the angular velocity, as expected from §3.2; this is well known in the vestibular literature (Benson, 1990). However, our results from this section suggest that when the forcing frequency is further increased (in humans, to about 100 Hz), the cupular deformation will be in phase with angular displacement. This sort of high frequency motion could be expected for example during impacts or equipment malfunction.